Main INDEX
Monthly INDEX
PREV
NEXT
User name Markowitz
Log entry time 00:02:19 on August 5,1999
Entry number 24294
keyword=shift summary
Started run with Hall in controlled access (for Euegene to do some Moller work) and accelerator trying to come back from a 2 hour injector tuning.
5 pm we hear we may be going to unpolarized beam for the next 2 days -- the runplan is according shifted.
We plan on doing the radiator test then taking gamma-d data -- but we note that the hadron arm momentum is presently at .658 not the 1.1424 listed in the runplan. When beam appears on the horizon we will call in Ron/Kuss to confirm we need to change. Radiator tests should not need the momentum change.
Javier has reset the epics stuff and the floor angles read from epics. However, HV is still not controlled from the FUI and must be done from the graphon.
1800 mcc calls to take hall A to beam permit. They expect us to have stable, unpolarized beam on target within the hour.
pulsed beam on target at 19:00. We have changed to a hadron angle of 61.5 and a momenta in the hadron arm of 1.1424 for the deuteron running. I.e., fields and momenta are set for the deuterium running (after the radiator check).
1940 beam in hall, 5 mikes. With no target and no radiator the count rate is unchanged from cosmics rates of 10 hz. putting in cryotarget and radiator.
with no target and 5 mikes and no radiator the rates were both 3 Hz e and h.
with no target and 5 mikes and 6% radiator the rate was 30 Hz electron and 10 Hz
hadron arm.
with deuterium target and 25 mikes and no radiator the rates were 22 KHz electron and 30 Hz hadron
with deuterium target and 25 mikes and the 6% radiator the rates were 40 khz in the electron arm and 300 hz in the hadron arm. Gilman concludes he is satisfied to run with the 6% radiator as it is below trip levels. (See entry 24272 in the halog.)
the raster is still 4 x 4 nominal
20:30 mcc calls. They lost control of one of the rf modules in the south linac, and hope to be back in about 15 minutes.
20:50 mcc calls. It will be about another 30 minutes they estimate.
21:45 mcc calls. It will be about 30 more minutes before we get beam.
22:05 mcc calls. Pulsed beam coming shortly.
22:15 beam back.
checked the spot size for the radiator tests and got about 5x3.5. Now adjustng raster size to the 0.8x1.0 that should give us the correct pattern of 2x2.
22:30 mcc calls. The calibration in hall C is causing problems and it will be 20 minutes before we have beam.
22:48 mcc calls. Beam will be up and ther raster has been changed to .8*1.0
run 720 is 6% radiator to check systematics of fpp
23:05 fpp chamber 4 trips off and is reset
23:12 beam gone
23:20 beam back
23:30 fp chamber 4 trips off and is reset
23:45 finish rate vs radiator thickness study and put in 6% radiator to take production data
23:59 fpp chamber 4 trips off and is reset
-----------------------------------------------
did some checks, started the first production run for photodisintegration of deuterium
but overall beam delivery was substandard.